Results for this submission Result 0 05 004 0.05 At least one of the answers above is NOT correct (1 point) An introduciory college course contains 45 students, 25 are Freshmen and 20 are Sophomores A student is chosen at random from those in the course, assume that each student is equalily likely Compute the following probabilties Give your answers in exact (fraction) form or round to at least three decimal places c The probabity of choosing a student who is not a Senior is 125 d The probablity of choosing a Junior is 1/20 sent to the LMS You have attempted this problem 1 tme You recerved a score or 0% for this attempt
Answer #1

a) P(freshman) = 25/45 or 0.556

b)P(Sophomore) = 20/45 or 0.444

c)P(not senior) = P(freshman) =  25/45 or 0.556

d) P(junior) = P(freshman) =  25/45 or 0.556